# Creative Problem Solving in Large Language and Vision Models – What Would it Take?

### v1

@@ Abstract
In this paper, we discuss approaches for integrating Computational Creativity (CC) with research in large language and vision models (LLVMs) to address a key limitation of these models, i.e., creative problem solving. We present preliminary experiments showing how CC principles can be applied to address this limitation through augmented prompting. With this work, we hope to foster discussions of Computational Creativity in the context of ML algorithms for creative problem solving in LLVMs. Our code is at: https://github.com/lnairGT/creative-problem-solving-LLMs
I Introduction
Creativity is “…the ability to come up with an idea which, relative to the pre-existing domain-space in one’s mind, one could not have had before. Whether any other person (or system) has already come up with it on an earlier occasion is irrelevant.[1], p.216. For artificial agents, Computational Creativity (CC) is a multi-disciplinary field (spanning Philosophy, Psychology, Neuroscience, and Computer Science) that seeks to develop computational methods capable of generating creative outcomes reminiscent of creative processes in humans [2]. Within CC, creative problem solving is a sub-area that requires an agent to discover – from its perspective – novel and previously unseen ways to accomplish a task. For example, in the absence of a ladle to scoop ingredients, an agent might creatively choose to substitute a bowl in place of the ladle. In this sense, creative problem solving encompasses creativity that is specifically task-oriented, as opposed to the generation of creative artefacts such as music or images.
While recent state-of-the-art large language and vision models (LLVMs) [^1] We use LLVM to denote the umbrella of large transformer models for natural language and vision tasks. Here, LLVMs include both LLMs and Vision-LMs or VLMs. have demonstrated competency in artistic endeavours [3, 4], creative problem solving continues to be a shortcoming of these models. For instance, in [5], the authors point out that “discontinuous tasks” that require a certain “Eureka” idea, i.e., creative problem solving, is currently a limitation of models like GPT-4. Similar observations have been made in follow up work showing that state-of-the-art LLMs inherently possess poor creative problem solving capabilities compared to humans [6]. Given this obvious limitation, ongoing research in Machine Learning should seek to address the gap between LLVMs and creative problem solving, to further enhance the intelligent capabilities of these models. As defined in prior work, “Intelligence is the ability to work and adapt to the environment with insufficient knowledge and resources.[7], p.10. Demonstrated in hallmark examples of human ingenuity, like the makeshift $CO_{2}$ filter built onboard the Apollo-13 [8], or the makeshift medical devices used to offset equipment shortages during COVID-19 [9], creative problem solving is especially important when dealing with resource-critical scenarios. However, such an exceptional degree of creative problem solving remains beyond the scope of LLVMs today.
@@ I Introduction
We believe that a discussion of Computational Creativity is essential to addressing this limitation. Machine Learning and Computational Creativity should be strongly integrated in research to enable effective creative problem solving in LLVMs and push the frontiers of their ingenuity. While [5] and [6] accurately point out creative problem solving as a shortcoming of state-of-the-art LLVMs, they do not expand their work to include the broader scope of Computational Creativity nor discuss how its principles can be applied to potentially alleviate this problem.
@@
This paper seeks to encourage the ML community to think about how LLVMs can be augmented with creative problem solving skills through a deeper discussion of Computational Creativity. To emphasize the applicability of principles from CC for creative problem solving in LLVMs, we discuss the seminal work of Margaret A. Boden from CC literature that introduces three forms of creativity, namely, “exploratory”, “combinational”, and “transformational[1]. While prior work has discussed the extension of Boden’s forms of creativity to creative problem solving in AI [2], their work does not include recent advances in LLVMs nor how Boden’s principles can be extended to specific approaches for these models.
@@
Ongoing discussions by leading ML experts like Dr. Shane Legg, co-founder of DeepMind, have suggested that “search” could help such models perform creative problem solving, quote, “… these foundational models are world models of a kind, and to do really creative problem solving, you need to start searching[10]. There has also been speculation that OpenAI’s $Q^{*}$ search (described as a “significant breakthrough” in popular media) could be targeting a similar approach [11, 12]. Interestingly, we note that “search” as described here, can be linked to Boden’s proposed “exploratory” approach (Section III-A1). However, we posit that “combinational” and “transformational” modes (Section III) should also be equally emphasized for achieving creative problem solving in LLVMs.
We begin by discussing the relation between creative problem solving and task planning, providing an overview of how LLVMs are used in task planning. We discuss the application of Boden’s three forms of creativity to enable creative problem solving in LLVMs. We also present preliminary experiments demonstrating the extension of transformational creativity to creative problem solving in LLVMs.
II Overview: LLVMs in typical task planning
@@
Creative problem solving is described as the process through which agents discover novel ways of accomplishing a task goal for a previously unsolvable task that can be computationally achieved through planning, learning, or hybrid methods [2]. In this section, we discuss how typical task planning is achieved with LLVMs. We divide the discussion into three subsections based on the level of task planning abstraction where LLVMs are applied: a) high-level task planning, b) low-level task planning, and c) hybrid task planning. While not exhaustive, our review is intended to provide a general insight into how LLVMs are used for task planning, to identify entry points for introducing creative problem solving capabilities.
II-A LLVMs for high-level task planning
@@
Approaches for high-level task planning often involve using LLVMs to identify high-level goals for accomplishing a task. Some approaches to task planning with LLMs often take a user input specifying the task, and generate high-level task plans for accomplishing it. These approaches often use LLMs as a form of “knowledge base”, to extract actionable task plans from the models via appropriate prompting [13], further iterating over the task plan with repeated calls to the LLM as needed [14].
@@
In the context of Reinforcement Learning (RL), prior work has focused on using LLMs to suggest high-level goals for an RL agent [15]. Dubbed as ELLMs (Exploring with LLMs), an RL agent provides its current state to an LLM via a prompt, and receives a goal suggestion from the LLM that is then used to shape the reward and the agent exploration. Further work has extended this approach to incorporate the use of experience memory [16]. Existing approaches have also used LLMs to generate directed acyclic graphs composed of sub-goal states to aid the exploration of an RL agent [17].
II-B LLVMs for low-level task planning
@@
Approaches for low-level task planning involve using LLMs to generate low-level code for performing a task. In contrast to high-level planning, where high-level goals and sub-goals are generated, these approaches use LLMs to directly generate low-level execution code via appropriate API calls [18]. Other approaches have also investigated the capacity of LLMs to generate task plans via a low-level planning language such as PDDL [19], including iterating over the generated plan descriptions in case of errors [20]. In terms of low-level planning using vision-based VLMs, prior work has introduced an approach that uses a diffusion model to generate robot trajectories conditioned on language and the current visual state of the robot [21].
II-C Hybrid high and low-level planning with LLVMs
@@
Hybrid approaches use LLVMs both for high-level goal generation as well as low-level planning. For instance, in [22], user inputs are passed as LLM prompts to generate high-level plans. The high-level plans are then converted to low-level plans for robot execution via LLMs specialized for coding. Other approaches have used a high-level LLM planner, a VLM perceiver, and a low-level LLM planner for re-planning with both visual and language inputs [23].
II-D Summary
@@
III Augmenting LLVM embedding spaces for creative problem solving
In this section, we discuss how principles from CC can be extended to LLVMs for creative problem solving. We begin with Boden’s definition of “conceptual spaces” as “[conceptual space] is the generative system that underlies the domain and defines a certain range of possibilities: chess moves, or molecular structures, or jazz melodies[24], p.18 and “… in short, any reasonably disciplined way of thinking[1], p.214. By this definition, the embedding space of an LLVM describes its conceptual space or “its way of thinking”. Some evidence for this also comes from existing work that introduces an approach for enabling LLMs to interpret continuous embedding spaces via natural language. Given an embedding vector representing an interpolation of different concepts, the model is able to interpret a text prompt in the context of the supplied embedding [25]. The embedding thus determines the model’s way of thinking. Hence, a discussion of enabling creative problem solving in LLVMs should target their embedding space. To this end, we explore two questions: a) how can LLVM embedding spaces be augmented to achieve creative problem solving, and b) what information should they be augmented with? Aligning with our original position, we show that CC literature can offer insights into these questions.
III-A How can LLVM embedding spaces be augmented?
@@
III-A1 Exploratory Creativity
Exploratory approaches involve exploration within the conceptual or equivalently, the embedding space of the model, and most closely relates to “search”. Note that the term “exploration” here differs from its usage in RL, instead referring to exploration through the model’s embedding space. Several existing approaches in the ML literature involve searching the output space of LLMs with the goal of improving the performance of these models. The “tree-of-thought” model generates a “tree” of next possible LLM outputs, and searches through the states via Breadth-first or Depth-first search to reach the desired goal state, often guided by heuristics [26]. Numerous other approaches have built upon a similar strategy, such as using Monte-Carlo Tree Search (MCTS) [27, 28], beam search [29] or integrating pruning to remove sub-par candidates [30].
Extension of exploratory creativity to LLVMs: An important point to note here is that these approaches involve searching exclusively within the output “solution space” of the LLMs rather than directly operating in the embedding space itself. In contrast to operating in the solution space of the LLM, exploratory approaches directly within the LLMs’ embedding space would not be limited by what the LLM can generate as output – “Some exploration merely shows us the nature of the relevant conceptual space that we had not explicitly noticed before[24], p.18. To effectively reveal the full extent of the conceptual space for creative problem solving, the approach should not be limited by the outputs the LLVM can generate. Rather, the generated (creative) outputs itself should be the result of heuristic or non-heuristic based search within the model’s embedding space. However, to the best of our knowledge current approaches have not focused on LLVMs from this perspective, and have also not applied search to embedding spaces of Vision-LMs. Regardless, exploratory approaches are still limited by the dimensions of the model’s embedding space. “To overcome a limitation in the conceptual space, one must change it in some way[24], p.18 - this leads us to combinational and transformational creativity.
III-A2 Combinational Creativity
Combinational approaches involve combining two concepts to create something new - “A novel combination of two familiar ideas is something which did not happen before.[1], p.213. We can broadly translate this to a function that takes in multiple concepts within an LLVM’s embedding space to output a novel concept.
One way of extending this definition to LLVMs involves applying cross-attention layers. The attention operation is defined as follows [31]:
@@ III-A2 Combinational Creativity
where, $Q$, $K$ and $V$ denote query, keys and values respectively, and $d_{k}$ denotes the dimensionality of the keys. Cross-attention involves passing $K$ and $V$ from a different model, e.g., in Flamingo [32], the keys and values represent visual input (from a separate vision encoder) and queries represent a language input. By applying cross attention in this manner, the embedding space of a model can be extended with capabilities of another model. In [33] the authors show that using cross-attention layers can help augment an anchor LLM with an augmenting LLM’s capabilities to perform a task that the anchor LLM was incapable of achieving before - hinting at some creative possibilities of this method.
@@
Other approaches in LLVMs, while using “combinations” in some way, do not conform to the notion of combinational creativity. This includes, for instance, approaches that perform arithmetic combination of LLM weights to enhance the model performance [34, 35]. Or approaches that combine image and text embeddings via concatenation [36] or a scaled dot product at the output [37]. While these approaches may be useful in imparting multi-modal capabilities, however, they do not lead to combinational creativity since the combination occurs external to the models as opposed to within the model’s embedding space.
@@
Extension of Combinational Creativity to LLVMs: The ML approaches described here involve combining embedding spaces across models. Existing approaches have not looked at combining concepts within the same model’s embedding space. The extension of combinational creativity to LLVMs is much more apparent in the sense of conceptual blending [38] for generation of creative artefacts, e.g., via blending of artistic styles. However, the extension of combinational creativity to creative problem solving is less obvious, and CC literature offers us further insights for making this connection. Typical conceptual blending corresponds to a form of “aesthetic combination”, whereas creative problem solving would benefit from “functional combinations” [39]. Functional combination combines the functions (as opposed to aesthetic) of two components, e.g., a coin combined with pliers could function as a makeshift screwdriver. The authors extend this framework to a combination of two nouns with a “base” noun (e.g., “pliers”) and “additive” noun (e.g., “coin”). An interesting possibility stems from this notion: Can a combination of embeddings of the same LLVM, corresponding to “base” and “additive” nouns (perhaps with some prior denoting the task), enable the LLVM to generate creative combinations of objects for solving a task? This question remains unexplored, and points to a potential research direction for LLVMs inspired by CC.
III-A3 Transformational Creativity

### v2

@@ Abstract
We advocate for a strong integration of Computational Creativity (CC) with research in large language and vision models (LLVMs) to address a key limitation of these models, i.e., creative problem solving. We present preliminary experiments showing how CC principles can be applied to address this limitation. Our goal is to foster discussions on creative problem solving in LLVMs and CC at prestigious ML venues.
Creative Problem Solving in Large Language and Vision Models – What Would it Take?
Lakshmi Nair Georgia Institute of Technology Atlanta, GA, USA                        Evana Gizzi Tufts University Medford, MA, USA                        Jivko Sinapov Tufts University Medford, MA, USA
1 Introduction
Creativity is “…the ability to come up with an idea which, relative to the pre-existing domain-space in one’s mind, one could not have had before. Whether any other person (or system) has already come up with it on an earlier occasion is irrelevant.Boden (1998), p.216. For artificial agents, Computational Creativity (CC) is a multi-disciplinary field (spanning Philosophy, Psychology, Neuroscience, and Computer Science) that seeks to develop computational methods capable of generating creative outcomes reminiscent of creative processes in humans Gizzi et al. (2022). Within CC, creative problem solving is a sub-area that requires an agent to discover – from its perspective – novel and previously unseen ways to accomplish a task. For example, in the absence of a ladle to scoop ingredients, an agent might creatively choose to substitute a bowl in place of the ladle. In this sense, creative problem solving encompasses creativity that is specifically task-oriented, as opposed to the generation of creative artifacts e.g., music or images.
@@ 1 Introduction
While recent state-of-the-art large language models (LLMs) and vision-language models (VLMs) have demonstrated competency in artistic endeavours Rombach et al. (2021); Copet et al. (2023), creative problem solving continues to be a shortcoming of these models (we use LLVM to denote the umbrella of both LLMs and VLMs). For instance, in Bubeck et al. (2023), the authors point out that “discontinuous tasks” that require a certain “Eureka” idea, i.e., creative problem solving, is currently a limitation of models like GPT-4. Similar observations have been made in follow up work showing that state-of-the-art LLMs inherently possess poor creative problem solving capabilities compared to humans Tian et al. (2023); Naeini et al. (2023). Given this obvious limitation, ongoing research in Machine Learning should seek to address the gap between LLVMs and creative problem solving, to further enhance the intelligent capabilities of these models. As defined in prior work, “Intelligence is the ability to work and adapt to the environment with insufficient knowledge and resources.Pennachin and Goertzel (2007), p.10. Demonstrated in hallmark examples of human ingenuity, like the makeshift $CO_{2}$ filter built onboard the Apollo-13 Cass (2005), or the makeshift medical devices used to offset equipment shortages during COVID-19 Turner et al. (2020), creative problem solving is especially important when dealing with resource-critical scenarios. Since humans may tend to “choke” under high pressure situations DeCaro et al. (2011) often limiting their CPS skills, autonomous agents equipped with LLVMs that have similar capabilities would be highly assistive and transformative to humans in high-stake environments. These include situations like rescue missions BBC (2012) or autonomous operation in human-inaccessible environments (e.g., space or underwater exploration) with limited resources Atkeson et al. (2018). However, the exceptional degree of creative problem solving necessary for such assistance remains beyond the scope of LLVMs today, limiting their intelligence (See Appx. B.1).
@@
We believe that a discussion of Computational Creativity is essential to addressing this limitation. It is our position that Machine Learning and Computational Creativity should be strongly integrated in research to enable effective creative problem solving in LLVMs and push the frontiers of their ingenuity.
2 Two Cultures Problem: Why does CC not receive a wider reception in ML?
Even though creative problem solving (CPS) is a shortcoming of existing LLVMs, Computational Creativity seldom finds its way into mainstream ML research. We believe this discrepancy aligns with the “two cultures” problem Hammond et al. (2013) (also corroborated in Van Heerden and Bas (2021); Lahikainen et al. (2024)), and is motivated by three aspects of CC literature as it relates to creative problem solving: a) the lack of a precise definition of CPS makes it challenging to identify how existing approaches in LLVMs are deficient in CPS skills; b) the somewhat “abstract” computational descriptions of CPS in Computational Creativity is challenging to connect to practical algorithms in LLVMs; and c) the lack of standardized benchmarks make it harder to evaluate LLVMs for CPS. In our discussions relating to a) in Section 3.1, b) in Section 4, and c) Section 5, we hope to address these gaps and encourage the ML community to think about how LLVMs can be augmented with creative problem solving skills through a deeper discussion of Computational Creativity.
@@ 2 Two Cultures Problem: Why does CC not receive a wider reception in ML?
To emphasize the applicability of principles from CC for creative problem solving in LLVMs, we discuss the seminal work of Margaret A. Boden from CC literature that introduces three forms of creativity, namely, “exploratory”, “combinational”, and “transformationalBoden (1998). Prior work has discussed the extension of Boden’s forms of creativity to creative problem solving in AI Gizzi et al. (2022), however, their work does not include recent advances in LLVMs nor how Boden’s principles can be extended to specific approaches for LLVMs.
@@
Ongoing discussions by leading ML experts like Dr. Shane Legg, co-founder of DeepMind, have suggested that “search” could help such models perform creative problem solving, quote, “… these foundational models are world models of a kind, and to do really creative problem solving, you need to start searchingPatel (2023). There has also been speculation that OpenAI’s $Q^{*}$ search (described as a “significant breakthrough” in popular media) could be targeting a similar approach Wang (2023); Anna Tong and Hu (2023). Interestingly, we note that “search” as described here, can be linked to Boden’s proposed “exploratory” approach (Section 4.1.1). However, in Section 4, we posit that “combinational” and “transformational” modes should be equally emphasized to achieve creative problem solving in LLVMs.
Although we choose to expand on Boden’s work as the focal point to drive our arguments in the main paper, it is not the only theory in CC that is relevant to this discussion. For completeness, we elaborate on additional CC theories and their applicability to creative problem solving in LLVMs in Appx. B.
3 From Task Planning to Creative Problem Solving
Creative problem solving can be broadly described as the process through which agents discover novel ways of accomplishing a task that, prior to the discovery, was unsolvable. Computationally, creative problem solving can be achieved through planning, learning, or hybrid approaches Gizzi et al. (2022). Following a review of the different definitions of creative problem solving that have been proposed (Appx. A), we believe the following most closely connects to existing formalisms in ML.
3.1 Definition of Creative Problem Solving
Gizzi et al. (2022) define the notion of a concept, as a state (of the environment and/or agent) or action. More generally, the authors denote $C_{X}$ as the set of all concepts relating to $X$ ($X$ denotes environment states $S$ or actions $A$). Hence, $C_{S}$ denotes the set of all environmental states, and $C_{A}$ denotes the set of agent actions. Formally, the authors state their definition as (Page 7, (Gizzi et al., 2022)):
Given an un-achievable goal due to an insufficient conceptual space, CPS refers to the process by which the agent discovers a new conceptual space $C^{\prime}_{X}\nsubseteq C_{X}$, such that $C^{\prime}_{X}=f(C_{X})$ is the result of applying some function $f$ on the current conceptual space, enabling the agent to solve the previously unsolvable task by using $C^{\prime}_{X}$.
As a simplified example, let us assume a robot that has a goal $G$ of transferring beans from a jar to a cooker: $G=${$in$(beans, cooker)}. Here, the initial state is defined as $C_{S}=${$in$(beans, jar), $hasContainability$(spoon)}. Let the actions be defined as $C_{A}=${$scoop$(beans, $X$, $loc_{s}$, $loc_{d}$)}, where, $X$ refers to an object that satisfies $hasContainability(\cdot)$ (e.g., spoon), to scoop beans from $loc_{s}$ to $loc_{d}$. If the robot has access to a spoon, the robot can use it to scoop the beans from the jar to the cooker. However, what if the robot did not have a spoon, but had a glass instead? By the definition of $C_{S}$, the agent is unaware that $hasContainability$(glass) is true, making the goal un-achievable. By our definition, creative problem solving is the process by which the agent uses some function $f(\cdot)$ to discover a new conceptual space: $f(C_{S})=C^{\prime}_{S}=C_{S}\mathop{\cup}${$hasContainability$ (glass)}. This would allow the agent to solve the previously unsolvable task by using the glass to scoop the beans instead.
Boden’s three forms of creativity denote three plausible functions for $f(C_{X})$. CPS arises when the agent uses what it knows, to discover something new and the newly discovered knowledge is applied to solve a previously impossible task. We revisit the notion of conceptual spaces in Section 3.
@@ 3.1 Definition of Creative Problem Solving
In the remainder of this section, we discuss how typical task planning is achieved with LLVMs. We divide the discussion into three subsections based on the level of task planning abstraction where LLVMs are applied: a) high-level task planning, b) low-level task planning, and c) hybrid task planning. While not exhaustive, our review is meant to offer a general insight into how LLVMs are used for task planning, to identify entry points for introducing creative problem solving capabilities.
3.2 LLVMs for high-level task planning
@@
Approaches for high-level task planning often involve using LLVMs to identify high-level goals for accomplishing a task. Some approaches to task planning with LLMs often take a user input specifying the task, and generate high-level task plans for accomplishing it. These approaches often use LLMs as a form of “knowledge base”, to extract actionable task plans from the models via appropriate prompting Huang et al. (2022), further iterating over the generated task plan with repeated calls to the LLM as needed Prasad et al. (2023).
@@
In the context of Reinforcement Learning (RL), prior work has focused on using LLMs to suggest high-level goals for an RL agent Du et al. (2023). Dubbed as ELLMs (Exploring with LLMs), an RL agent provides its current state to an LLM via a prompt, and receives a goal suggestion from the LLM that is then used to shape the reward and the agent exploration. Further work has extended this approach to incorporate the use of experience memory Zhang et al. (2023a). Existing approaches have also used LLMs to generate directed acyclic graphs composed of sub-goal states to aid the exploration of an RL agent Shukla et al. (2023).
3.3 LLVMs for low-level task planning
@@
Approaches for low-level task planning involve using LLMs to generate low-level code for performing a task. In contrast to high-level planning, where high-level goals and sub-goals are generated, these approaches use LLMs to directly generate low-level execution code via appropriate API calls Liang et al. (2023). Other approaches have also investigated the capacity of LLMs to generate task plans via a low-level planning language such as PDDL Silver et al. (2023), including iterating over the generated plan descriptions in case of errors Guan et al. (2023). In terms of low-level planning using VLMs, prior work has introduced an approach that uses a diffusion model to generate robot trajectories conditioned on language and the current visual state of the robot Chen et al. (2023).
3.4 Hybrid high and low-level planning with LLVMs
@@
Hybrid approaches use LLVMs both for high-level goal generation as well as low-level planning. For instance, in Li et al. (2023), user inputs are passed as LLM prompts to generate high-level plans. The high-level plans are then converted to low-level plans for robot execution via LLMs specialized for coding. Other approaches have used a high-level LLM planner, a VLM perceiver, and a low-level LLM planner for re-planning with both visual and language inputs Skreta et al. (2024).
3.5 Summary
@@
4 Augmenting LLVM embedding spaces for creative problem solving
In this section, we discuss how principles from CC can be extended to LLVMs for creative problem solving. We begin with Boden’s definition of “conceptual spaces” as “[conceptual space] is the generative system that underlies the domain and defines a certain range of possibilities: chess moves, or molecular structures, or jazz melodiesBoden (2005), p.18 and “… in short, any reasonably disciplined way of thinkingBoden (1998), p.214. By this definition, the embedding space of an LLVM describes its conceptual space or “its way of thinking”. Some evidence for this also comes from existing work that introduces an approach for enabling LLMs to interpret continuous embedding spaces via natural language. Given an embedding vector representing an interpolation of different concepts, the model is able to interpret a text prompt in the context of the supplied embedding Tennenholtz et al. (2023). The embedding thus determines the model’s way of thinking. Hence, a discussion of enabling creative problem solving in LLVMs should target their embedding space. To this end, we explore two questions: a) how can LLVM embedding spaces be augmented to achieve creative problem solving, and b) what information should they be augmented with? Aligning with our original position, we show that CC literature can offer insights into these questions.
4.1 How can LLVM embedding spaces be augmented?
@@
4.1.1 Exploratory Creativity
Exploratory approaches involve exploration within the conceptual or equivalently, the embedding space of the model, and most closely relates to “search”. Note that the term “exploration” here differs from its usage in RL, instead referring to exploration through the model’s embedding space. Several existing approaches in the ML literature involve searching the output space of LLMs with the goal of improving the performance of these models. The “tree-of-thought” model generates a “tree” of next possible LLM outputs, and searches through the states via Breadth-first or Depth-first search to reach the desired goal state, often guided by heuristics Yao et al. (2023). Numerous other approaches have built upon a similar strategy, such as using Monte-Carlo Tree Search (MCTS) Zhou et al. (2023); Feng et al. (2023), beam search Zhang et al. (2023b) or integrating pruning to remove sub-par candidates Golovneva et al. (2023).
Extension of exploratory creativity to LLVMs: An important point to note here is that these approaches involve searching exclusively within the output “solution space” of the LLMs rather than directly operating in the embedding space itself. In contrast to operating in the solution space of the LLM, exploratory approaches directly within the LLMs’ embedding space would not be limited by what the LLM can generate as output – “Some exploration merely shows us the nature of the relevant conceptual space that we had not explicitly noticed beforeBoden (2005), p.18. To effectively reveal the full extent of the conceptual space for creative problem solving, the approach should not be limited by the outputs the LLVM can generate. Rather, the generated (creative) outputs itself should be the result of heuristic or non-heuristic based search within the model’s embedding space. However, to the best of our knowledge current approaches have not focused on LLVMs from this perspective, and have also not applied search to embedding spaces of Vision-LMs. Regardless, exploratory approaches are still limited by the dimensions of the model’s embedding space. “To overcome a limitation in the conceptual space, one must change it in some wayBoden (2005), p.18 - this leads us to combinational and transformational creativity.
4.1.2 Combinational Creativity
Combinational approaches involve combining two concepts to create something new - “A novel combination of two familiar ideas is something which did not happen before.Boden (1998), p.213. We can broadly translate this to a function that takes in multiple concepts within an LLVM’s embedding space to output a novel concept.
One way of extending this definition to LLVMs involves applying cross-attention layers. The attention operation is defined as Vaswani et al. (2017):
@@ 4.1.2 Combinational Creativity
where, $Q$, $K$ and $V$ denote query, keys and values respectively, and $d_{k}$ denotes the dimensionality of the keys. Cross-attention involves passing $K$ and $V$ from a different model, e.g., in Flamingo Alayrac et al. (2022), the keys and values represent visual input (from a separate vision encoder) and queries represent a language input. By applying cross attention in this manner, the embedding space of a model can be extended with capabilities of another model. In Bansal et al. (2024) the authors show that using cross-attention layers can help augment an anchor LLM with an augmenting LLM’s capabilities to perform a task that the anchor LLM was incapable of achieving before - hinting at some creative possibilities of this method.
@@
Other approaches in LLVMs, while using “combinations” in some way, do not conform to the notion of combinational creativity. This includes, for instance, approaches that perform arithmetic combination of LLM weights to enhance the model performance Matena and Raffel (2022); Ilharco et al. (2022). Or approaches that combine image and text embeddings via concatenation Kim et al. (2021) or a scaled dot product at the output Radford et al. (2021). While these approaches may be useful in imparting multi-modal capabilities, however, they do not lead to combinational creativity since the combination occurs external to the models as opposed to within the model’s embedding space.
@@
Extension of Combinational Creativity to LLVMs: The ML approaches described here involve combining embedding spaces across models. Existing approaches have not looked at combining concepts within the same model’s embedding space. The extension of combinational creativity to LLVMs is much more apparent in the sense of conceptual blending Fauconnier and Turner (2003) for generation of creative artifacts, e.g., via blending of artistic styles. However, the extension of combinational creativity to creative problem solving is less obvious, and CC literature offers us further insights for making this connection. Typical conceptual blending corresponds to a form of “aesthetic combination”, whereas creative problem solving would benefit from “functional combinations” Chen et al. (2018). Functional combination combines the functions (as opposed to aesthetic) of two components, e.g., a coin combined with pliers could function as a makeshift screwdriver. The authors extend this framework to a combination of two nouns with a “base” noun (e.g., “pliers”) and “additive” noun (e.g., “coin”). An interesting possibility stems from this notion: Can a combination of embeddings of the same LLVM, corresponding to “base” and “additive” nouns (perhaps with some prior denoting the task), enable the LLVM to generate creative combinations of objects for solving a task? This question remains unexplored, and points to a potential research direction for LLVMs inspired by CC.
4.1.3 Transformational Creativity